# The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals

### v1

@@ 4.1 RQ1: Understanding of PGs
Skepticism on the privacy statement and limits of privacy-preserving guarantees. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that "no privacy technology is foolproof". Similarly, in the for-profit scenario, P45 wrote: "I don’t trust that the privacy-preserving tech would work." More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: "Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane!". Others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that "the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked", while P124 noted that "corporate data breaches are very common", and P80 echoed this sentiment, saying: "I believe the [intentions] will be good, but data can be hacked". We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: "They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen".
Doubt on the recipient entity’s motivation to employ PETs. 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities "will do what is profitable and not much more than that". Similarly, P119 stated "(…) because it is a for-profit organization. I expect them to cut corners" and P32 wrote that "for-profit organization have low standard of morality". However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: "Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans".
4.2 RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see section 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?". As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
Statement Shown Overall For-Profit Non-Profit
PG(1) - Anonymization 0.68 0.63 0.73
+ AG(1) - Expert audit 0.63 0.59 0.68
+ AG(2) - Self audit 0.70 0.67 0.74
PG(2) - Access control 0.68 0.67 0.68
+ AG(1) - Expert audit 0.69 0.59 0.78
+ AG(2) - Self audit 0.58 0.61 0.56
PG(3) - Data expiration 0.73 0.67 0.80
+ AG(1) - Expert audit 0.69 0.67 0.72
+ AG(2) - Self audit 0.68 0.70 0.65
PG(4) - Purpose restriction 0.72 0.53 0.89
+ AG(1) - Expert audit 0.71 0.71 0.70
+ AG(2) - Self audit 0.72 0.72 0.72
Control 0.65 0.60 0.68
TABLE V: Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario.(e.g., the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.)
@@ 4.2 RQ2: Willingness to donate health data
Analysis Set Up.Analysis Set UpAnalysis Set Up.
Table V summarizes the respondents’ willingness to donate their health data to a recipient entity. The 1st column (“overall") reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table VI.
@@
   Education 0.75 [0.33, 1.63] 0.466
   Age 0.68 [0.33, 1.39] 0.290
   Gender 0.80 [0.40, 1.61] 0.535
   Tech Background 1.33 [0.57, 3.15] 0.512
   Egocentrism 0.41 [0.20, 0.84] 0.016
   Donation History 9.47 [2.22, 54.68] 0.005
@@
   Education 0.37 [0.15, 0.86] 0.025
   Age 1.28 [0.61, 2.70] 0.516
   Gender 0.83 [0.38, 1.78] 0.630
   Tech Background 0.39 [0.16, 0.89] 0.026
   Egocentrism 1.79 [0.85, 3.96] 0.136
   Donation History 4.11 [1.23, 17.58] 0.034
@@
TABLE VI: Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
Figure 6: Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (305). $N=456$.
@@
Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
There was no direct relationship between showing a privacy statement and respondents’ willingness to donate. However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 6, we observe high alignment among the three explored constructs, where 234 (51.3%) respondents gave all positive responses and 70 (15.4%) respondents gave all negative responses. Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
Privacy statements and expectations are not the only reason for donation intentions however. While participants were generally less willing to donate to for-profit entities (OR=0.66, $p=0.038$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities. This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with egocentric relationships to the donation subject are less willing to donate to a for-profit entity (OR = 0.41, $p=0.016$). This is potentially related to their existing (negative) experiences with for-profit entities in healthcare. For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information. P244 also mentioned that their information has been leaked before by a clinician. An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 "just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
Respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.37, $p=0.025$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
Prior donors are more willing to donate to both entities. Respondents who have donated before are more willing to donate to both entities, with the OR value of 9.47 ($p=0.005$) and 4.11 ($p=0.034$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes." and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection."
@@
Qualitative Analysis of Donation Intention Via Open-Answer Responses.Qualitative Analysis of Donation Intention Via Open-Answer ResponsesQualitative Analysis of Donation Intention Via Open-Answer Responses.
One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that "there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines". Similarly, P97 stated: "even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in".
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned donating because "my dad died when his heart failed on him, and I’d have rather that not had happened", and P233 noted that they "would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did".
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that "I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled".
Respondents expressed negative sentiments toward for-profit entities collecting data for research. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: "I have a negative connotation with for profit organization" and "I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me." Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: "For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount."
Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations. For instance, P122 mentions: "it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit". Then, P282 notes: "it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have."
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that "there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future." Others cited previous data leak experiences as to why they would not donate, as P290 said "my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others." Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that "regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.". Another reason was that the participant was concerned that their medical data included too much or too specific information. P155 highlighted this concern: "I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not"; as did P229: "It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me." Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., "I don’t have mayor illnesses or nothing to hide so Im ok with that") or because they already share information with other entities (e.g., "Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different").
5 Discussion
Analyzing the effect of privacy statements on users’ perceptions, trust, and willingness to donate in the context of a survey is particularly useful because the survey mimics the actual process of consenting that recipient entities use when requesting data. These entities typically provide an introduction, present privacy guarantees, and ask for consent to donate data, much like our survey.Our study reveals several key findings about privacy expectations, their influence on willingness to donate health data, and the implications for both researchers and regulators.
@@ 5 Discussion
Pre-existing Privacy Expectations.Pre-existing Privacy ExpectationsPre-existing Privacy Expectations. We find that people have strong pre-existing privacy expectations even when no specific privacy statements are shown. This expectation is particularly high for non-profit entities, which are perceived as more ethical, thus prioritizing public good over profit. In contrast, for-profit entities are seen as driven by financial gain, leading to skepticism on their commitment to protecting user privacy. Non-profit entities foster greater trust and confidence in their data handling practices to the point that respondents often expect more stringent privacy measures, such as purpose use restriction. However, such measures may not be technically enforced or commonly practiced; indeed, many health organizations employ broad (not purpose restricted) consent [58]. This gap between expectations and reality can lead to privacy expectation violations, which are a significant source of privacy concerns according to the theory of contextual integrity [77].
@@
Impact of Privacy Statements.Impact of Privacy StatementsImpact of Privacy Statements. Privacy statements play a role similar to advertising, such as those used by companies like Apple [79]. For for-profit entities, where baseline privacy expectations are lower, privacy statements raise these expectations. (In contrast, for non-profits, the impact is minimal because the expectations are already high.) Thus, we argue that privacy expectations of for-profits can be effectively influenced by transparently implementing and communicating about PETs.

### v2

@@ 4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
Skepticism on the privacy statement and limits of privacy-preserving guarantees..Skepticism on the privacy statement and limits of privacy-preserving guarantees.Skepticism on the privacy statement and limits of privacy-preserving guarantees.. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that ”no privacy technology is foolproof”. Similarly, in the for-profit scenario, P45 wrote: ”I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: ”Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane!”. Some others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that ”the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that ”corporate data breaches are very common”, and P80 echoed this sentiment, saying: ”I believe the [intentions] will be good, but data can be hacked”. We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: ”They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen”.
[…]
Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs.. 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities ”will do what is profitable and not much more than that”. Similarly, P119 stated ”(…) because it is a for-profit organization. I expect them to cut corners” and P32 wrote that ”for-profit organization have low standard of morality”. However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: ”Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans”.
4.2. RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see § 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?”. As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
4.2.1. Analysis Set Up
In Figure 9, we summarize the respondents’ willingness to donate their health data to a recipient entity. The 1st column (“overall”) reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table 3.
@@ 4.2.1. Analysis Set Up
4.2.2. Response Distribution and Statistical Analysis Results
Figure 9. Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
Figure 10. Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310). $N=454$.
@@ 4.2.2. Response Distribution and Statistical Analysis Results
   Education 0.73 [0.43, 1.26] 0.264
   Age 1.00 [0.98, 1.01] 0.765
   Gender 1.30 [0.80, 2.12] 0.304
   Tech Background 1.57 [0.88, 2.80] 0.127
   Egocentricity 0.69 [0.42, 1.14] 0.144
   Donation History 3.25 [1.31, 8.01] 0.011
@@
   Education 0.46 [0.26, 0.80] 0.007
   Age 1.00 [0.98, 1.02] 0.897
   Gender 1.16 [0.69, 1.97] 0.560
   Tech Background 0.71 [0.40, 1.26] 0.237
   Egocentricity 1.36 [0.83, 2.23] 0.217
   Donation History 2.25 [1.06, 4.75] 0.033
@@
Table 3. Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
There was no direct relationship between showing a privacy statement and respondents’ willingness to donate. However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 10, we observe high alignment among the three explored constructs, where 236 (52.0%) respondents gave all positive responses and 63 (13.9%) respondents gave all negative responses. Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
Privacy statements and expectations are not the only reason for donation intentions however. While participants were generally less willing to donate to for-profit entities (OR=0.38, $p=\textless 0.001$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities. This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p=0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
Finally, we found that prior donors are more willing to donate to both entities. Respondents who have donated before are more willing to donate to both entities, with the OR value of 3.25 ($p=0.011$) and 2.25 ($p=0.033$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes.” and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection”.
4.2.3. Qualitative Analysis of Donation Intention Via Open-Answer Responses
One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that ”there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines.” Similarly, P97 stated: ”even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in.”
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned donating because ”my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they ”would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did.”
However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations. For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information. P244 also mentioned that their information has been leaked before by a clinician.
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that ”I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.”
An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 ”just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
Respondents expressed negative sentiments toward for-profit entities collecting data for research. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: ”I have a negative connotation with for profit organization” and ”I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: ”For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount.”
Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations. For instance, P122 mentions: ”it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit.” Then, P282 notes: ”it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that ”there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said ”my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that ”regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”.
Another reason was that the participant was concerned that their medical data included too much or too specific information. P155 highlighted this concern: ”I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: ”It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., ”I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., ”Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
5. Concluding Discussion
Like our survey, when soliciting data donations, recipient entities usually provide an introduction motivating their need to collect the individual’s data, outline privacy guarantees, and seek consent for data donation. Privacy statements function similarly to advertising (O’Flaherty, 2022) by shaping and influencing consumers’ expectations of privacy (Schaub et al., 2015). Yet, these statements do not always have impact.
Prior work has highlighted the ways in which such statements can fail due to being vague or inaccurate. Our work highlights how the impact of privacy statements is moderated by people’s pre-existing privacy expectations.
Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements. We observe that these expectations vary strongly based on the profit model of the entity. For for-profit entities, where baseline privacy expectations are lower, strong privacy statements can effectively elevate these expectations among users. In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities.
@@ 5. Concluding Discussion
The Halo Effect.The Halo EffectThe Halo Effect. Participants qualitative responses suggest that their strongly positive privacy expectations for non-profit entities are rooted in arguments of morality: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit. This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates (Nisbett and Wilson, 1977; de Bruin Cardoso et al., 2024). While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
These positive expectations do not always align with reality. Privacy safeguards are not consistently enforced or widely practiced; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures (Kaye et al., 2015). This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity (Nissenbaum, 2009). Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry (Levine, 2022). Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., (Burton et al., 2015)).
@@
Privacy Washing.Privacy WashingPrivacy Washing. Conversely, our participants’ had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect). As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
This means that for-profit entities have the opportunity to gain “competitive advantage” (Schaub et al., 2015) through influencing and enhancing user privacy expectations via transparently implementing and effectively communicating about PETs. However, this effect also posses a vulnerability to ”privacy washing” (MainWP, 2023; Cirucci, 2024), in which organizations make vague or exaggerated claims about PETs to falsely raise privacy expectations. Such deceptive practices can mislead individuals into sharing their data, erode trust, and discourage future data donations if individuals later feel deceived.
@@
The Role of Auditing.The Role of AuditingThe Role of Auditing. In theory, auditing could offer verification that combats both the halo effect and privacy washing by grounding assumed or stated privacy guarantees in technical reality. Unfortunately, we find that auditing has a limited effect on elevating privacy expectations among participants, except in scenarios involving purpose use restrictions by for-profit entities. Participants understand that while PETs and auditing processes are in place, they are not infallible, and auditing does not substantially mitigate the practical risks associated with data breaches or unauthorized access.
This perception shows a disconnect between the perspectives of experts and participants, as shown in (Dechand et al., 2019; Oates et al., 2018; Wu and Zappala, 2018). Experts often regard auditing as a fundamental assurance mechanism that reinforces trust and compliance. In contrast, participants are more concerned with the overall protection of their data, regardless of specific technical implementations or compliance processes. As one respondent noted, ”They can guarantee privacy all they want; things still get hacked.” True privacy is doubted to exist (Lerner et al., 2022).
@@
Building Trust.Building TrustBuilding Trust. Our respondent’s prior experiences with data leakage often informed this fatalistic perspective on privacy. Technologies like multi-party computation (MPC), although not yet scalable, have the potential to minimize the impact of a breach by specifically distributing sensitive information across different entities to ensure that no single entity can reconstruct the information on its own. However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved. Existing skepticism about data security means that any failure of these systems could severely damage trust, potentially more so than if simpler, well-understood methods were compromised.